All Systems Operational
Voice / Messaging API's Operational
Inbound Voice Operational
Outbound Voice Operational
Inbound SMS Operational
Outbound SMS Operational
Transcription ? Operational
api.tropo.com (Tokens) ? Operational
api.sl.tropo.com Operational
Provisioning / Hosting Operational
api.tropo.com (Provisioning) ? Operational
Webhosting Operational
ftp.tropo.com ? Operational
Tropo.com Website ? Operational
Connect Operational
api.connect.tropo.com ? Operational
connect.tropo.com ? Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Scheduled Maintenance
Hosted maintenance next week Apr 28, 21:00 - Apr 29, 01:00 EDT
This week Tropo Operations will be performing the following maintenance tasks in our hosted environment:

* Updates to load balancers for api.tropo.com to implement security changes and improve general performance and reliability
* Update expiring SSL certificate for Phono.com
* Update Token API's w/ new build to address an edge case where a single token launch could result in two sessions being launched
* Update internal vendor certificate used for www.tropo.com used for fraud prevention

If there are any questions please let us know, as our team is most certainly standing by to address them ( support@tropo.com )
Posted on Apr 17, 12:30 EDT
Past Incidents
Apr 28, 2015

No incidents reported today.

Apr 27, 2015

No incidents reported.

Apr 26, 2015

No incidents reported.

Apr 25, 2015

No incidents reported.

Apr 24, 2015

No incidents reported.

Apr 23, 2015

No incidents reported.

Apr 22, 2015
Resolved - We have identified an issue with a subset of our speech recognition servers. The problematic servers have been removed from routing and at this point tests are green. We will continue to monitor the situation and if there are any questions please let us know ( support@tropo.com ).
Apr 22, 20:51 EDT
Investigating - We are seeing some reports of user experiencing failures when using non-english ASR languages in our production environment. This issue is NOT impacting English ( en-us ) speech recognition. Our Operations engineers are currently investigating these reports.
Apr 22, 19:37 EDT
Resolved - This incident has been resolved.
Apr 22, 19:38 EDT
Monitoring - Approximately 2 hours ago Tropo Operations was given the green light to fully route back to our primary data center and this work has been completed. All monitoring and checks are passing and all primary services are deemed fully restored at this time. If there are any questions please don't hesitate to let us know ( support@tropo.com ).
Apr 20, 22:42 EDT
Update - All calls and SMS should be working without issue at this point, though datacenter problems still persist and likely will throughout much of the day. We will continue to monitor and update as we have more information, or things change.
Apr 20, 06:55 EDT
Update - We have routed away and voice is working, as is outbound SMS; inbound SMS appears to still have a potential problem and we are continuing to work on it. There was approximately a one hour window where some voice calls may have failed (2 a.m. - 3 a.m. EST) prior to the successful route away. Please contact support@tropo.com with questions, and we will continue to provide updates as we have them.
Apr 20, 04:01 EDT
Identified - Our datacenter provider is reporting a water leak in their building; we are routing away to keep production traffic functioning, but the development environment will be down for the time being until we are able to route back. We will update as we have more information.
Apr 20, 02:38 EDT
Apr 21, 2015

No incidents reported.

Apr 19, 2015

No incidents reported.

Apr 18, 2015

No incidents reported.

Apr 17, 2015

No incidents reported.

Apr 16, 2015

No incidents reported.

Apr 15, 2015

No incidents reported.

Apr 14, 2015
Resolved - This incident has been resolved.
Apr 14, 15:23 EDT
Monitoring - Further investigation indicates the issue was load related, and only occurred on a single production runtime, and only impacted a single active call. The issue was also very brief and cleared almost immediately. We suspect it was an out of control users application which triggered the high load. We are continuing to monitor the cluster at this time.
Apr 14, 13:13 EDT
Investigating - Tropo monitoring is showing some potential issues with the replication conference data in our shared production environment. We are taking action but there appears to have been at least 1 user impacted already by this event.
Apr 14, 12:47 EDT